OpenStudy (anonymous):

Make 20 with (4,4,5,5)????

OpenStudy (anonymous):

4x5+20x(5-1)=20x1+20

OpenStudy (anonymous):

4x5=20*

OpenStudy (anonymous):

5*4/5-4 or 5*4*(5-4) or .........

Parth (parthkohli):

\[4 \times 5\times(5 - 4) \]

hartnn (hartnn):

or \[\sqrt{4*5*4*5}\]

OpenStudy (anonymous):

Dynamic programming algorithm.

Parth (parthkohli):

Or\[4^2 - 4^2 + 5^2 - 5\]

Parth (parthkohli):

So that is in order too! ^

OpenStudy (anonymous):

nice 2x

Parth (parthkohli):

In a similar spirit,\[(4 - 4) + 5^2 - 5\]

hartnn (hartnn):

4^2+4 +5-5 but i guess, we used 2 in 'square'